Find the profitability index (PI) for the following series of future cash flows, assuming the company’s cost of capital is 5.84 percent. The initial outlay is $485,058. Year 1: $196,161 Year 2: $147,113 Year 3: $129,342 Year 4: $175,446 Year 5: $162,596 Round the answer to two decimal places.
Answer #1

Profitability index = PV of cash flows/ Initial cost

Pv of cash flows = CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 196161/1.0584^1+ 147113/1.0584^2+ 129342/1.0584^3 + 175446/1.0584^4+ 162596/1.0584^5

=687989.05

Profitability index = PV of cash flows/ Initial cost

=687989.05/ 485058

=1.42